OpenStudy (anonymous):

Becky invested $32,000 at 6% interest compounded annually. What will be her total balance at the end of 2 years?

OpenStudy (anonymous):

f(x) = 32000(1.06)^2 I believe plug 2 in for x!

OpenStudy (anonymous):

1.1236?

OpenStudy (anonymous):

My bad i think its f(x) = 32000(1.06)2 32000 x 2.12 = 67840

OpenStudy (anonymous):

No, your answer is incorrect. I just checked that, sorry D: Look how I did.. A = P ( 1 + r ) ^ t = 32,000 * ( 1 + 0.06 ) ^ 2 = 32,000 * ( 1.06 ) ^ 2 = 32,000 * 1.1236 = 35955.2 My final answer is 35955.2 but that's incorrect too. ;-;

OpenStudy (anonymous):

I see...so its not an exponent nor a whole number?

OpenStudy (anonymous):

I don't know that yet

OpenStudy (anonymous):

@robtobey @radar @ParthKohli @phi where did we go wrong?

OpenStudy (anonymous):

@Jadeishere

OpenStudy (jadeishere):

$32,000 at 6% interest equals to $1920. In two years Becky will have $35840.

OpenStudy (anonymous):

http://prntscr.com/3pwliw nvrm. I got the answer anyway :P My answer was correct. I only had to add a 0 at the end. Like this: 35955.20 Thanks for helping though ;)

OpenStudy (jadeishere):

>.<